# The coherent structure of the energy cascade in isotropic turbulence

### v3

@@ Results
$\Pi^{S}_{nl}$ $=$ $-\frac{1}{12}\int_{0}^{r^{2}}\left(\overline{\overline{S}_{ki}^{\theta}\overline{S}_{jk}^{\theta}}^{\phi}-\overline{\overline{S}_{ki}^{\theta}}^{\phi}\overline{\overline{S}_{jk}^{\theta}}^{\phi}\right)\overline{S}_{ij}^{r}\mathrm{d}\theta^{2},$ (8)
@@
$\Pi^{C}_{nl}$ $=$ $-\frac{1}{12}\int_{0}^{r^{2}}\left(\overline{\overline{S}_{ki}^{\theta}\overline{\Omega}_{jk}^{\theta}}^{\phi}-\overline{\overline{S}_{ki}^{\theta}}^{\phi}\overline{\overline{\Omega}_{jk}^{\theta}}^{\phi}\right)\overline{S}_{ij}^{r}\mathrm{d}\theta^{2},$ (10)
@@
where $\phi=\sqrt{r^{2}-\theta^{2}}$ and $\Omega_{ij}$ is the rate-of-rotation tensor.
@@
The terms $\Pi_{l}^{S}$ and $\Pi_{l}^{\omega}$ (where $l$ denotes ‘local’) are the energy transfer due to local-in-scale interactions of $\overline{S}_{ij}^{r}$ with either the rate-of-strain tensor $\overline{S}_{ij}^{r}$ (i.e., strain self-amplification) or the vorticity vector $\overline{\omega}_{i}^{r}$ (i.e, vortex stretching), respectively. The terms $\Pi^{S}_{nl}$ and $\Pi^{\omega}_{nl}$ (where $nl$ denotes ‘non-local’) are analogous to $\Pi_{l}^{S}$ and $\Pi_{l}^{\omega}$ but correspond to nonlocal-in-scale interactions. Despite the non-local nature of $\Pi^{S}_{nl}$ and $\Pi^{\omega}_{nl}$, it has been shown that they still represent interactions occurring close in scale [21, 23, 48, 49, 25]. The remaining term, $\Pi_{nl}^{C}$ is also nonlocal in scale and represents energy transfer by the resolved strain-rate tensor acting on the product of strain-rate and vorticity.
@@
The conditional-averaged values for the total energy transfer ($\{\Pi\}$) and its individual contributions ($\{\Pi_{l}^{S}\}$, $\{\Pi_{l}^{\omega}\}$, $\{\Pi_{nl}^{S}\}$, $\{\Pi_{nl}^{\omega}\}$, and $\{\Pi_{nl}^{C}\}$) are calculated following the ensemble average procedure from §Conditional-averaged flow fields. For ease of visualization, the values are plotted in figure 10 along the $\hat{x}$ coordinate, but a similar picture emerges along the other directions, $\hat{y}$ and $\hat{z}$. The results reveal that roughly 85% of the overall energy transfer is primarily attributed to strain-rate self-amplification ($\{\Pi_{l}^{S}\}+\{\Pi_{nl}^{S}\}$), most of which is local (approximately 70%), whereas the contribution of vortex stretching ($\{\Pi_{l}^{\omega}\}+\{\Pi_{nl}^{\omega}\}$) accounts for less than 15%. The dominant role of the local strain-rate self-amplification is consistent with the saddle point topology of the conditional-averaged velocity field from figure 6 and the staggered arrangement between enstrophy and energy transfer from figure 7. Interestingly, the contribution of $\{\Pi_{l}^{\omega}\}$ is essentially zero, implying that vortex stretching acts only from larger-scale strain to slightly smaller-scale vorticity via the term $\{\Pi_{nl}^{\omega}\}$. This phenomenon has been observed in previous studies [31, 32, 36, 37]. The contribution of $\Pi_{nl}^{C}$ is also negligible, at least for the range of filter widths considered. This agrees with the space-time average values of $\Pi_{nl}^{C}$ reported in the literature [48, 49]. Although some sensitivities can be appreciated across the filter widths (figure 10a) and Reynolds numbers (figure 10b), the conclusions remain robust across the cases investigated.
@@ Conclusions
The main focus of this work has been on the coherent structure and physical mechanisms involved in the energy cascade. Nonetheless, our results can inform decisions about subgrid-scale (SGS) model developments for LES, especially when the intent is to faithfully represent the processes involved in the cascade. In such situations, the coherent structure of the energy cascade must be consistent with the predominant role of strain-self amplification over vortex stretching. The high strain-rate region created between two distinct zones of elevated enstrophy, as reported here, can serve as a benchmark to evaluate the physical fidelity of SGS models in those contexts. This level of detail is probably not required for SGS models that are aimed at capturing only the general characteristics of the flow, such as mean velocity profiles or mean Reynolds stresses. Another significant implication for SGS modeling stems from the observation that, although 70% of the energy transfer ($\approx\{\Pi_{l}^{S}\}$) is due to local interactions (and thus resolvable by LES), the remaining 30% arises from nonlocal interactions. These latter cannot be resolved by the LES grid and therefore need to be modeled, introducing the challenge of the closure problem.
@@ Acknowledgements
A.L.-D. conceived the idea, D.P. conducted the data post-processing and developed the code, D.P. and A.L.-D. analysed the results, All authors reviewed the manuscript.

### v4

@@ Results
$\Pi^{S}_{nl}$ $=$ $-\frac{1}{12}\int_{0}^{r^{2}}\left(\overline{\overline{S}_{ik}^{\theta}\overline{S}_{jk}^{\theta}}^{\phi}-\overline{\overline{S}_{ik}^{\theta}}^{\phi}\overline{\overline{S}_{jk}^{\theta}}^{\phi}\right)\overline{S}_{ij}^{r}\mathrm{d}\theta^{2},$ (10)
@@
$\Pi^{C}_{nl}$ $=$ $-\frac{1}{12}\int_{0}^{r^{2}}\left(\overline{\overline{S}_{ik}^{\theta}\overline{\Omega}_{jk}^{\theta}}^{\phi}+\overline{\overline{S}_{ik}^{\theta}}^{\phi}\overline{\overline{\Omega}_{jk}^{\theta}}^{\phi}\right)\overline{S}_{ij}^{r}\mathrm{d}\theta^{2},$ (12)
@@
where $\phi=\sqrt{r^{2}-\theta^{2}}$ and $\Omega_{ij}=(\partial u_{i}/\partial x_{j}-\partial u_{j}/\partial x_{i})/2$ is the rate-of-rotation tensor.
@@
The terms $\Pi_{l}^{S}$ and $\Pi_{l}^{\omega}$ (where $l$ denotes ‘local’) are the energy transfer due to local-in-scale interactions of $\overline{S}_{ij}^{r}$ with either the rate-of-strain tensor $\overline{S}_{ij}^{r}$ (i.e., strain self-amplification) or the vorticity vector $\overline{\omega}_{i}^{r}$ (i.e, vortex stretching), respectively. The terms $\Pi^{S}_{nl}$ and $\Pi^{\omega}_{nl}$ (where $nl$ denotes ‘non-local’) are analogous to $\Pi_{l}^{S}$ and $\Pi_{l}^{\omega}$ but correspond to non-local-in-scale interactions. Despite the non-local nature of $\Pi^{S}_{nl}$ and $\Pi^{\omega}_{nl}$, it has been shown that they still represent interactions occurring close in scale [21, 23, 49, 50, 25]. The remaining term, $\Pi_{nl}^{C}$ is also non-local in scale and represents energy transfer by the resolved strain-rate tensor acting on the product of strain-rate and vorticity.
@@
The conditional-averaged values for the total energy transfer ($\{\Pi\}$) and its individual contributions ($\{\Pi_{l}^{S}\}$, $\{\Pi_{l}^{\omega}\}$, $\{\Pi_{nl}^{S}\}$, $\{\Pi_{nl}^{\omega}\}$, and $\{\Pi_{nl}^{C}\}$) are calculated following the ensemble average procedure from figure 5. For ease of visualization, the values are plotted in figure 10 along the $\hat{x}$ coordinate, but a similar picture emerges along the other directions, $\hat{y}$ and $\hat{z}$. The results reveal that roughly 85% of the overall energy transfer is primarily attributed to strain-rate self-amplification ($\{\Pi_{l}^{S}\}+\{\Pi_{nl}^{S}\}$), most of which is local (approximately 70%), whereas the contribution of vortex stretching ($\{\Pi_{l}^{\omega}\}+\{\Pi_{nl}^{\omega}\}$) accounts for less than 15%. The dominant role of the local strain-rate self-amplification is consistent with the saddle point topology of the conditional-averaged velocity field from figure 6 and the staggered arrangement between enstrophy and energy transfer from figure 7. Interestingly, the contribution of $\{\Pi_{l}^{\omega}\}$ is essentially zero, implying that vortex stretching acts only from larger-scale strain to slightly smaller-scale vorticity via the term $\{\Pi_{nl}^{\omega}\}$. This phenomenon has been observed in previous studies [32, 33, 37, 38]. The contribution of $\Pi_{nl}^{C}$ is also negligible, at least for the range of filter widths considered. This agrees with the space-time average values of $\Pi_{nl}^{C}$ reported in the literature [49, 50]. Although some sensitivities can be appreciated across the filter widths (figure 10a) and Reynolds numbers (figure 10b), the conclusions remain robust across the cases investigated.
@@ Conclusions
The main focus of this work has been on the coherent structure and physical mechanisms involved in the energy cascade. Nonetheless, our results can inform decisions about subgrid-scale (SGS) model developments for LES, especially when the intent is to faithfully represent the processes involved in the cascade. In such situations, the coherent structure of the energy cascade must be consistent with the predominant role of strain-self amplification over vortex stretching. The high strain-rate region created between two distinct zones of elevated enstrophy, as reported here, can serve as a benchmark to evaluate the physical fidelity of SGS models in those contexts. This level of detail is probably not required for SGS models that are aimed at capturing only the general characteristics of the flow, such as mean velocity profiles or mean Reynolds stresses. Another significant implication for SGS modeling stems from the observation that, although 70% of the energy transfer ($\approx\{\Pi_{l}^{S}\}$) is due to local interactions (and thus resolvable by LES), the remaining 30% arises from non-local interactions. These latter cannot be resolved by the LES grid and therefore need to be modeled, introducing the challenge of the closure problem.
@@
Data availability
The datasets generated and analyzed during the current study are available in the Jimenez’s Group repository, https://torroja.dmt.upm.es/turbdata/
@@ Acknowledgements
A.L.-D. conceived the idea, D.P. conducted the data post-processing and developed the code, D.P. and A.L.-D. analyzed the results. All authors reviewed the manuscript.